• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Boeglin

    Log entry time 16:01:42 on August 9,2006

    Entry number 178685

    keyword=Shift Summary

    8:00 continue data on LD2 with radiator in
    10:05 beam stopped. Hall will be put in controlled access
    10:55 MCC called, Hall ready for access
    Spectrometer back at its place
    Electronic work finished: redundant helicity measurements and checks added
    13:50 Hall Closed, ask for beam permit, prepare for Bull's Eye scan
    14:30 start harp scans
    15:25 Harp scanner B fails revert to normal pointing run


    Run, Events, Total, Setup
    2406/20815, 1.95, 36, LD2 rad. in
    2407/20815, 1.95, 37.9, "
    2408/20815, 1.95, 39.85, "
    2409/20815, 1.95, 41.8, "
    2410/20816, 1.95, 43.75, "
    2411/20817, 1.95, 45.7, "
    2412/20818, 1.95, 47.65, "

    Harp scans:
    Position, Harp run, Coda runs
    (0,0), 2247, 2425/20820
    A(0.9,0.9)
    B(0.,0.) 2248, 2426/20821 (end of run at wrong position)

    A(-.6,.2)
    B(-2.,-2.) 2249, 2427/20822

    Mini-Pointing runs
    2429 pointing at 5uA carbon
    2430 pointing at 20uA
    2431 pointing, Dummy 4cm